# The Identification Power of

### v1

@@ 3.2 Characterization of $\Theta_{IC}$ for Combined Data
We next consider characterizing the identified set $\Theta_{IC}$ under the combined data. Given that the experimental data identify $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$, we begin by examining what additional information can be obtained by incorporating the observational data. The following lemma provides an answer to this question.
Lemma 3.1 (Identification of $F_{Y_{1}S}^{\ast}$ and $F_{Y_{0}S}^{\ast}$).
@@
Lemma 3.1 implies that by combining experimental and observational data, we can identify the joint distributions $\left(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\right)$ of each potential outcome, the self-selection variable, and covariates. Hence, the combined data provide richer information than the experimental data alone, which identify only $\left(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast}\right)$.
Specifically, $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)$ for $d\neq s$ is a counterfactual distribution and cannot be identified from either the experimental or observational data alone. However, by combining both data sources, it can be identified via the following decomposition: For $d\neq s$,
$F_{Y_{d}|X}^{\ast}(\cdot|x)=\mathbb{P}(S=d|X=x)\cdot F_{Y_{d}|SX}^{\ast}(\cdot|d,x)+\mathbb{P}(S=s|X=x)\cdot F_{Y_{d}|SX}^{\ast}(\cdot|s,x),$
where $F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)$ is identified from the experimental data, while $\mathbb{P}(S=\cdot\mid X=x)$ and $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid d,x)$ are identified from the observational data (see the proof for details). [^8] Long et al. (2008) also show a related result: the identification of the ATE $\mathbb{E}[Y_{1}-Y_{0}\mid S=s]$ for each self-selection group $s\in\{0,1\}$ by combining data from random assignment and self-selection sources.
Let $C^{\ast}(\cdot,\cdot|s,x)$ denote the true conditional copula given $S=s$ and $X=x$, that is,
@@
$\theta_{o}$ $=\mathbb{E}_{F_{Y_{1}Y_{0}}^{\ast}}\left[\psi(Y_{1},Y_{0})\right]=\mathbb{E}_{F_{SX}^{\ast}}\left[\mathbb{E}_{F_{Y_{1}Y_{0}|SX}^{\ast}}\left[\psi(Y_{1},Y_{0})\middle|S,X\right]\right]$
@@
Although the true conditional copula $C^{\ast}(\cdot,\cdot\mid s,x)$ is unknown, by replacing it with all possible copulas $C(\cdot,\cdot|s,x)\in\mathcal{C}$, we can construct the identified set for $\theta_{o}$ based on the knowledge of $\bigl{(}F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr{)}$ as follows:
@@
where, as implied by Sklar’s theorem, for any $C\in\mathcal{C}$, the function $C\big{(}F_{Y_{1}|SX}^{\ast}(y_{1}\mid s,x),F_{Y_{0}|SX}^{\ast}(y_{0}\mid s,x)\mid s,x\big{)}$ defines a valid conditional joint CDF of $(Y_{1},Y_{0})$ given $S=s$ and $X=x$, whose marginal CDFs correspond to $F_{Y_{1}|SX}^{\ast}(\cdot\mid s,x)$ and $F_{Y_{0}|SX}^{\ast}(\cdot\mid s,x)$.
We now turn to the question of whether $\widetilde{\Theta}_{IC}$ coincides with the identified set $\Theta_{IC}$ under the combined data. This question can be rephrased as whether the self-selection joint distributions $\bigl{(}F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr{)}$ fully characterize the identified set $\Theta_{IC}$. This is a nontrivial question, as combining experimental and observational data might yield additional identifying information beyond what is captured by these joint distributions.
The following theorem provides an affirmative answer to this question by showing that the identified set $\Theta_{IC}$ is indeed fully characterized by the joint distributions $\bigl{(}F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr{)}$.
@@
The identified set $\Theta_{IC}$ under the combined data is equivalent to the identified $\widetilde{\Theta}_{IC}$ based on the knowledge of $\bigl{(}F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr{)}$; that is, $\Theta_{IC}=\widetilde{\Theta}_{IC}$.
To summarize, the identified set $\Theta_{I}$ under the experimental data is fully characterized by the joint distributions $(F_{Y_{0}X}^{\ast},F_{Y_{1}X}^{\ast})$ (Proposition 3.1), while the identified set $\Theta_{IC}$ under the combined data is fully characterized by the self-selection joint distributions $(F_{Y_{0}SX}^{\ast},F_{Y_{1}SX}^{\ast})$ (Theorem 3.2). This comparison suggests that combining experimental and observational data brings additional identifying power through the joint distribution of each potential outcome and the self-selection variable. The following section investigates this point in detail.
@@
The comparison between $\widetilde{\Theta}_{I}$ and $\widetilde{\Theta}_{IC}$ shows that if self-selection $S$ is independent of the potential outcomes $(Y_{1},Y_{0})$ conditional on $X$, then $\Theta_{I}$ and $\Theta_{IC}$ coincide; that is, combining the experimental and observational data does not tighten the identified set. This implies that if selection-on-observables, $(Y_{1},Y_{0})\perp\!\!\!\perp D\mid X$, holds in the observational data, the combination of the two data sources provides no additional identifying power for $\theta_{o}$. In contrast to many identification analyses where selection-on-observables in the observational data is desirable, self-selection on unobservables is preferable in the present context.
@@ 4 Bounds Analysis
In this section, we explicitly characterize the sharp bounds for the DTE parameter $\theta_{o}$ under both experimental and combined data. The analysis builds on Fan et al. (2017), who study the identifying power of covariates for distributional treatment effects. Extending their framework, we examine the identifying power of the self-selection variable $S$, which, through Theorem 3.2, reflects the identifying power of the combined data.
Following Fan et al. (2017), we consider two classes of DTE parameters, where $\psi$ is specified as (i) a super-modular function and (ii) a $\varphi$-indicator function. For each case, we derive analytical expressions for the sharp bounds on $\theta_{o}$ under both experimental and combined data, and we characterize the necessary and sufficient conditions under which combining experimental and observational data tightens the identified set (i.e., $\Theta_{IC}\subset\Theta_{I}$).
4.1 Characterization of the Identified Sets for Super-modular Functions and the Identification Power of Combined Data
We begin by considering the DTR parameters represented by super-modular and sub-modular functions.
@@
(i) A function $\psi(\cdot,\cdot)$ is said to be super-modular if, for all $y_{1}\leq y_{1}^{\prime}$ and $y_{0}\leq y_{0}^{\prime}$, $\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})\geq 0$. It is sub-modular if $\psi(\cdot,\cdot)$ is super-modular. (ii) A function $\psi(\cdot,\cdot)$ is called strict super-modular if, for all $y_{1}<y_{1}^{\prime}$ and $y_{0}<y_{0}^{\prime}$, $\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})>0$. It is strict sub-modular if $\psi(\cdot,\cdot)$ is strict super-modular.
Cambanis et al. (1976) provide many examples of super-modular and sub-modular functions; see also Fan et al. (2017). The functions $\psi(\cdot,\cdot)$ in Examples 2.3, 2.4, and 2.5 are either super-modular or sub-modular and therefore fall within this framework. In particular, the function $\psi$ in Example 2.5 is strict super-modular. Fan et al. (2017) also present further examples of strict super-modular and sub-modular functions.
@@
where $M(u,v)\equiv\max(u+v-1,0)$ and $W(u,v)\equiv\min(u,v)$ are the Fréchet–Hoeffding lower and upper bounds, respectively, on the joint distribution.
When $\psi$ is super-modular, Theorem 3.2 of Fan et al. (2017) shows that, under certain regularity conditions, the identified set $\widetilde{\Theta}_{I}$ based on $(F_{Y_{1}X}^{\ast},F_{Y_{1}X}^{\ast})$ is the interval $\left[\theta^{L},\theta^{U}\right]$, where
@@
with $F_{Y_{d}|X}^{\ast,-1}(u|x)\equiv\inf\left\{y:F_{Y_{d}|X}^{\ast,-1}(y|x)\geq u\right\}$ denoting the quintile function of $y_{d}$ conditional on $X=x$ for $d=0,1$.
Therefore, by Proposition 3.1, the identified set $\Theta_{I}$ from the experimental data is given by the interval $\left[\theta^{L},\theta^{U}\right]$. Fan et al. (2017) also provide inference methods based on the analytical expressions (5) and (6) of the sharp bounds.
@@
where the difference from $F^{\ast,(-)}(y_{1},y_{0})$ and $F^{\ast,(+)}(y_{1},y_{0})$ lies in the inclusion of the self-selection variable $S$ in the conditioning set.
Then, applying Theorem 3.2 of Fan et al. (2017), we obtain that when $\psi$ is super-modular and certain regularity conditions hold, the identified set $\widetilde{\Theta}_{IC}$ based on $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$ is the interval $\left[\theta_{L},\theta_{U}\right]$, where
@@
with $F_{Y_{d}|SX}^{\ast,-1}(u|s,x)\equiv\inf\{y:F_{Y_{d}|SX}^{\ast,-1}(y|s,x)\geq u\}$ denoting the quintile function of $y_{d}$ conditional on $S=s$ and $X=x$ for $d=0,1$.
Therefore, by Theorem 3.2, the identified set $\Theta_{IC}$ from the combined data is also given by $\left[\theta_{L},\theta_{U}\right]$; that is, $\Theta_{IC}=\left[\theta_{L},\theta_{U}\right]$.
We formalize these results in the following proposition.
@@
    Suppose that $\theta^{L}$ and $\theta^{U}$ exist (possibly infinite), and that either of the following conditions holds: (a) $\psi(y_{1},y_{0})$ is symmetric, and both $\mathbb{E}[\psi(Y_{1},Y_{1})]$ and $\mathbb{E}[\psi(Y_{0},Y_{0})]$ are finite; (b) there exist some constants $\bar{y}_{0}$ and $\bar{y}_{1}$ such that $\mathbb{E}[\psi(Y_{1},\bar{y}_{0})]$ and $\mathbb{E}[\psi(\bar{y}_{1},Y_{0})]$ are finite, and at least one of $\theta^{L}$ and $\theta^{U}$ is finite. Then, $\Theta_{I}=\left[\theta^{L},\theta^{U}\right]$.
@@
Proposition 4.2 (ii) is novel in that it characterizes the sharp bounds for $\theta_{o}$ under the combination of experimental and observational data when $\psi$ is super-modular. Together with Lemma 3.1, this result enables the computation of these bounds using the combined data via the analytical expressions in equations (7) and (8), along with equations (3) and (4).It also provides a basis for constructing inference procedures.
The key difference between the sharp bounds $\theta^{L}$ ($\theta^{U}$) under the experimental data and $\theta_{L}$ ($\theta_{U}$) under the combined data lies in the inclusion of the self-selection variable $S$ in the conditioning sets (see equations (7) and (8)). This inclusion implies that self-selection can offer additional identifying power for $\theta_{o}$ when experimental data are combined with observational data. In particular, the variable $S$ is expected to encode information about the dependence structure between $Y_{1}$ and $Y_{0}$—for example, as suggested by the Roy selection model—and can thereby tighten the identified set.
For strict modular functions $\psi$, Theorem 4.2 below establishes necessary and sufficient condition for $\Theta_{IC}=\Theta_{I}$, or equivalently, for $\Theta_{IC}$ to be a proper subset of $\Theta_{I}$, meaning that the combination of experimental and observational data strictly tightens the identified set.
@@
Let $\psi(y_{1},y_{0})$ be a strict super-modular and right continuous function. Suppose that the four expectations in (5)–(6) and (7)–(8) exist (even if infinite valued), and that either conditions (a) and (b) in Proposition 4.1 (i) holds. Then, $\Theta_{IC}=\Theta_{I}$ if and only if the following two conditions hold for $\psi_{c}$-almost all $(y_{1},y_{0})$: [^9] If $\psi(\cdot,\cdot)$ is super-modular and right continuous, it uniquely determines a non-negative measure $\psi_{c}$ on the Borel subsets of $\mathbb{R}^{2}$ such that for all $y_{1}\leq y_{1}^{\prime}$ and $y_{0}\leq y_{0}^{\prime}$, $\psi_{c}((y_{1},y_{1}^{\prime}]\times(y_{0},y_{0}^{\prime}])=\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})$. See Cambanis et al. (1976) and Rachev and Rüschendorf (2006).
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\right)\in\{0,1\}\mbox{\ and\ }$ (9)
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\right)\in\{0,1\}.$ (10)
The conditions in equations (9) and (10) are restrictive and unlikely to hold when the self-selection variable $S$ depends on the potential outcomes $Y_{1}$ and $Y_{0}$. In such cases, the conditional distributions $F_{Y_{1}|S,X}^{\ast}$ and $F_{Y_{0}|S,X}^{\ast}$ are likely to violate these conditions. Therefore, when self-selection is outcome-dependent, combining experimental and observational data is likely to tighten the identified set for $\theta_{o}$. A notable exception arises when self-selection is independent of the potential outcomes conditional on covariates XX—that is, when selection-on-observables holds. In this case, conditions (9) and (10) are satisfied, and the data combination offers no additional identifying power.
4.2 Characterization of the Identified Sets for $\varphi$-indicator Functions and the Identification Power of Data Combination

### v4

@@ 3.2 Characterization of $\Theta_{IC}$ for Combined Data
We next seek to characterize the identified set $\Theta_{IC}$ under the combined data. Given that the experimental data identify $(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast})$, our starting point is to examine what additional information can be obtained by incorporating the observational data. The following lemma addresses this question.
Lemma 3.2 (Identification of $F_{Y_{1}SX}^{\ast}$ and $F_{Y_{0}SX}^{\ast}$).
@@
Lemma 3.2 shows that combining experimental and observational data enables identification of the joint distributions $\left(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\right)$ of each potential outcome, self-selection variable, and covariates. Therefore, the combined data yield richer information than the experimental data alone, which identify only $\left(F_{Y_{1}X}^{\ast},F_{Y_{0}X}^{\ast}\right)$.
In particular, $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid s,x)$ for $d\neq s$ is a counterfactual distribution and cannot be identified from either the experimental or observational data alone. Combining both data sources, however, makes it identifiable through the following decomposition: For $d\neq s$,
$F_{Y_{d}|X}^{\ast}(\cdot|x)=\mathbb{P}(S=d|X=x)\cdot F_{Y_{d}|SX}^{\ast}(\cdot|d,x)+\mathbb{P}(S=s|X=x)\cdot F_{Y_{d}|SX}^{\ast}(\cdot|s,x).$
Here, $F_{Y_{d}\mid X}^{\ast}(\cdot\mid x)$ is identified from the experimental data, while $\mathbb{P}(S=\cdot\mid X=x)$ and $F_{Y_{d}\mid SX}^{\ast}(\cdot\mid d,x)$ are identified from the observational data (see the proof for details). [^7] Long et al. (2008) also show a related result: the ATE $\mathbb{E}[Y_{1}-Y_{0}\mid S=s]$ for each self-selection group $s\in\{0,1\}$ can be identified by combining data from random assignment and self-selection sources.
Let $C^{\ast}(\cdot,\cdot|s,x)$ denote the true conditional copula given $S=s$ and $X=x$; that is,
@@
$\theta_{o}$ $=\mathbb{E}_{F_{Y_{1}Y_{0}}^{\ast}}\left[\psi(Y_{1},Y_{0})\right]$
@@
Although the true conditional copula $C^{\ast}(\cdot,\cdot\mid s,x)$ is unknown, replacing it with all possible copulas $C(\cdot,\cdot|s,x)\in\mathcal{C}$ yields the identified set for $\theta_{o}$ based on $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$:
@@
By Sklar’s theorem, for any $C\in\mathcal{C}$, the function $C\big(F_{Y_{1}|SX}^{\ast}(y_{1}\mid s,x),F_{Y_{0}|SX}^{\ast}(y_{0}\mid s,x)\mid s,x\big)$ defines a valid conditional joint CDF of $(Y_{1},Y_{0})$ given $S=s$ and $X=x$, with marginals $F_{Y_{1}|SX}^{\ast}(\cdot\mid s,x)$ and $F_{Y_{0}|SX}^{\ast}(\cdot\mid s,x)$.
We now ask whether $\widetilde{\Theta}_{IC}$ coincides with the identified set $\Theta_{IC}$ under the combined data. This question can be rephrased as whether $\Theta_{IC}$ is fully characterized by the self-selection joint distributions $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$. This question is nontrivial because combining experimental and observational data might yield additional identifying information beyond these joint distributions.
The following theorem establishes an affirmative answer to this question: the identified set $\Theta_{IC}$ is fully characterized by $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$.
@@
The identified set $\Theta_{IC}$ under the combined data is equivalent to the identified set $\widetilde{\Theta}_{IC}$ based on the distributions $\bigl(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast}\bigr)$; that is, $\Theta_{IC}=\widetilde{\Theta}_{IC}$.
In summary, the identified set $\Theta_{I}$ under the experimental data is fully characterized by $(F_{Y_{0}X}^{\ast},F_{Y_{1}X}^{\ast})$ (Proposition 3.1), whereas the identified set $\Theta_{IC}$ under the combined data is fully characterized by the self-selection joint distributions $(F_{Y_{0}SX}^{\ast},F_{Y_{1}SX}^{\ast})$ (Theorem 3.3). This comparison suggests that combining experimental and observational data yields additional identifying power through the dependence between potential outcomes $(Y_{1},Y_{0})$ and self-selection $S$. The following section examines this point in detail.
@@
The comparison between $\widetilde{\Theta}_{I}$ and $\widetilde{\Theta}_{IC}$ reveals that if self-selection $S$ is independent of the potential outcomes $(Y_{1},Y_{0})$ conditional on $X$, then $\Theta_{I}$ and $\Theta_{IC}$ coincide. This implies that when selection-on-observables, $(Y_{1},Y_{0})\perp\!\!\!\perp D\mid X$, holds in the observational data, the combination of the two data sources provides no additional identifying power for $\theta_{o}$. In contrast to much of the identification literature, where selection-on-observables is desirable in observational data, identification benefits from selection-on-unobservables in the present context.
@@ 4 Bounds Analysis
In this section, we derive sharp bounds for the DTE parameter $\theta_{o}$ under both experimental and combined data. Our analysis builds on Fan et al. (2017), who study the identifying power of covariates for DTEs. Extending their framework, we examine the identifying power of the self-selection variable $S$, which, as shown in Theorem 3.3, captures the additional identifying power of the combined data.
Following Fan et al. (2017), we consider two classes of DTE parameters, where $\psi$ is specified as (i) a super-modular function and (ii) a $\varphi$-indicator function. For each case, we derive closed-form expressions for the sharp bounds on $\theta_{o}$ under both experimental and combined data. We further establish, for each case, necessary and sufficient conditions under which combining the two data sources strictly tightens the identified set (i.e., $\Theta_{IC}\subset\Theta_{I}$).
4.1 Identified Sets for Super-modular Functions and the Identification Power of Combined Data
We begin with DTR parameters represented by super-modular and sub-modular functions.
@@ 4.1 Identified Sets for Super-modular Functions and the Identification Power of Combined Data
(i) A function $\psi(\cdot,\cdot)$ is super-modular if, for all $y_{1}\leq y_{1}^{\prime}$ and $y_{0}\leq y_{0}^{\prime}$, $\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})\geq 0$. It is sub-modular if $\psi(\cdot,\cdot)$ is super-modular. (ii) A function $\psi(\cdot,\cdot)$ is strict super-modular if, for all $y_{1}<y_{1}^{\prime}$ and $y_{0}<y_{0}^{\prime}$, $\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})>0$. It is strict sub-modular if $\psi(\cdot,\cdot)$ is strict super-modular.
The functions $\psi(\cdot,\cdot)$ in Examples 2.3, 2.4, and 2.5 are either super-modular or sub-modular and thus fall within this framework. In particular, the function $\psi$ in Example 2.5 is strict super-modular. Cambanis et al. (1976) provide many other examples of super-modular and sub-modular functions; see also Fan et al. (2017) for further examples of strict super-modular and sub-modular functions.
@@
where $M(u,v)\equiv\max(u+v-1,0)$ and $W(u,v)\equiv\min(u,v)$ are the Fréchet–Hoeffding lower and upper bounds, respectively, for a bivariate distribution with marginals $(u,v)$.
When $\psi$ is super-modular, Theorem 3.2 of Fan et al. (2017) shows that under certain regularity conditions, the identified set $\widetilde{\Theta}_{I}$ based on $(F_{Y_{1}X}^{\ast},F_{Y_{1}X}^{\ast})$ is the interval $\left[\theta^{L},\theta^{U}\right]$, where
@@
with $F_{Y_{d}|X}^{\ast,-1}(u|x)\equiv\inf\bigl\{y:F_{Y_{d}|X}^{\ast,-1}(y|x)\geq u\bigr\}$ denoting the quintile function of $Y_{d}$ conditional on X=xX=x. [^8] These analytical expressions highlight the identification power of the covariates $X$, as shown by Fan et al. (2017).
Therefore, by Proposition 3.1, the identified set $\Theta_{I}$ under the experimental data is given by the interval $\left[\theta^{L},\theta^{U}\right]$. Fan et al. (2017) also propose inference methods based on the analytical expressions (6) and (7) for the sharp bounds.
@@
where $F_{(\star)}^{\ast}(y_{1},y_{0})$, for $\star\in\{-,+\}$, differs from $F^{\ast,(\star)}(y_{1},y_{0})$ by the inclusion of the self-selection variable $S$.
Then, by applying Theorem 3.2 of Fan et al. (2017), we obtain that when $\psi$ is super-modular and certain regularity conditions hold, the identified set $\widetilde{\Theta}_{IC}$ based on $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$ is the interval $\left[\theta_{L},\theta_{U}\right]$, where
@@
with $F_{Y_{d}|SX}^{\ast,-1}(u|s,x)\equiv\inf\{y:F_{Y_{d}|SX}^{\ast,-1}(y|s,x)\geq u\}$ denoting the quintile function of $Y_{d}$ conditional on $S=s$ and $X=x$.
Therefore, by Theorem 3.3, the identified set $\Theta_{IC}$ under the combined data is given by the interval $\left[\theta_{L},\theta_{U}\right]$. We formalize these results in the following proposition.
@@
    Suppose that $\theta^{L}$ and $\theta^{U}$ exist (possibly infinite), and that either of the following conditions holds: (a) $\psi(y_{1},y_{0})$ is symmetric and both $\mathbb{E}[\psi(Y_{1},Y_{1})]$ and $\mathbb{E}[\psi(Y_{0},Y_{0})]$ are finite; (b) there exist some constants $\bar{y}_{0}$ and $\bar{y}_{1}$ such that $\mathbb{E}[\psi(Y_{1},\bar{y}_{0})]$ and $\mathbb{E}[\psi(\bar{y}_{1},Y_{0})]$ are finite, and at least one of $\theta^{L}$ and $\theta^{U}$ is finite. Then, $\Theta_{I}=\left[\theta^{L},\theta^{U}\right]$.
@@
Proposition 4.1(ii) is our novel contribution, deriving the sharp bounds for $\theta_{o}$ under the combined experimental and observational data when $\psi$ is super-modular. Together with Lemma 3.2, this result enables computation of the sharp bounds via equations (8) and (9), in conjunction with (4) and (5). It also provides a basis for constructing inference procedures. [^9] Pointwise valid confidence sets for $\theta_{o}$ can be constructed by extending the inference procedure of Fan et al. (2017, Appendix B) to settings with nonparametric estimators of $(F_{Y_{1}SX}^{\ast},F_{Y_{0}SX}^{\ast})$, which can be obtained using Lemma 3.2.
The key difference between the sharp bounds $\theta^{L}$ ($\theta^{U}$) and $\theta_{L}$ ($\theta_{U}$) under the two data scenarios lies in the inclusion of the self-selection variable $S$ in the conditioning sets (see equations (8) and (9)). This inclusion implies that self-selection can offer additional identifying power for $\theta_{o}$ when experimental data are combined with observational data. In particular, $S$ may carry information about the dependence structure between $Y_{1}$ and $Y_{0}$ (for example, as in the Roy selection model), thereby tightening the identified set.
To see this formally, note that $\Theta_{I}$ is the set of values of $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$ as $F$ ranges over all joint CDFs $F_{Y_{1}Y_{0}}$ with fixed marginals $F^{\ast,(-)}$ and $F^{\ast,(+)}$ satisfying $F^{\ast,(-)}(y_{1},y_{0})\leq F_{Y_{1}Y_{0}}(y_{1},y_{0})\leq F^{\ast,(+)}(y_{1},y_{0})$, whereas $\Theta_{IC}$ is the set of values of $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$ as $F$ ranges over all joint CDFs $F_{Y_{1}Y_{0}}$ with fixed marginals $F_{(-)}^{\ast}$ and $F_{(+)}^{\ast}$ satisfying $F_{(-)}^{\ast}(y_{1},y_{0})\leq F_{Y_{1}Y_{0}}(y_{1},y_{0})\leq F_{(+)}^{\ast}(y_{1},y_{0})$. By Jensen’s inequality, it follows that
$F^{\ast,(-)}(y_{1},y_{0})$ $=\mathbb{E}\left[\max\left\{\mathbb{E}\left[F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1\middle|X\right],0\right\}\right]$
$=F_{(-)}^{\ast}(y_{1},y_{0}),$
and similarly $F^{\ast,(+)}(y_{1},y_{0})\geq F_{(+)}^{\ast}(y_{1},y_{0})$. Since $F^{\ast,(-)}(\cdot,\cdot)\leq F_{(-)}^{\ast}(\cdot,\cdot)$ and $F_{(+)}^{\ast}(\cdot,\cdot)\leq F^{\ast,(+)}(\cdot,\cdot)$, we obtain $\theta^{L}\leq\theta_{L}$ and $\theta_{U}\leq\theta^{U}$, which implies that the data combination potentially shrinks the identified set.
For strict modular functions $\psi$, Theorem 4.2 below establishes the necessary and sufficient condition under which $\Theta_{IC}=\Theta_{I}$; otherwise, combining the two data sources yields a strictly smaller identified set (i.e., $\Theta_{IC}\subsetneq\Theta_{I}$).
@@
Let $\psi(y_{1},y_{0})$ be a strict super-modular and right-continuous function. Suppose that the four expectations in equations (6)–(9) exist (even if infinite valued), and that either conditions (a) and (b) in Proposition 4.1 (i) hold. Then $\Theta_{IC}=\Theta_{I}$ if and only if the following two conditions hold for $\psi_{c}$-almost all $(y_{1},y_{0})$: [^10] If $\psi(\cdot,\cdot)$ is super-modular and right continuous, it uniquely determines a non-negative measure $\psi_{c}$ on the Borel subsets of $\mathbb{R}^{2}$ such that for all $y_{1}\leq y_{1}^{\prime}$ and $y_{0}\leq y_{0}^{\prime}$, $\psi_{c}((y_{1},y_{1}^{\prime}]\times(y_{0},y_{0}^{\prime}])=\psi(y_{1},y_{0})+\psi(y_{1}^{\prime},y_{0}^{\prime})-\psi(y_{1},y_{0}^{\prime})-\psi(y_{1}^{\prime},y_{0})$. See Cambanis et al. (1976) and Rachev and Rüschendorf (2006).
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\middle|X\right)\in\{0,1\}\text{ a.s. }\mbox{\ and\ }$ (10)
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\middle|X\right)\in\{0,1\}\text{ a.s.}$ (11)
The conditions in equations (10) and (11) require that, conditional on $X$, the relative ordering of the conditional distributions $F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)$ and $F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)$ is degenerate in the sense that it does not vary with the self-selection variable $S$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Such conditions are highly restrictive and unlikely to hold when self-selection $S$ depends on the potential outcomes $Y_{1}$ and $Y_{0}$. In such cases, the conditional distributions $F_{Y_{1}|S,X}^{\ast}$ and $F_{Y_{0}|S,X}^{\ast}$ typically vary with $S$ in a way that violates these conditions. Therefore, when self-selection is outcome-dependent, combining experimental and observational data is likely to tighten the identified set for $\theta_{o}$.
A notable exception arises when self-selection is independent of the potential outcomes conditional on $X$ (i.e., when selection-on-observables holds). In this case, equations (10) and (11) are satisfied, and combining the two data sources provides no additional identifying power.
4.2 Identified Sets for $\varphi$-indicator Functions and the Identification Power of Data Combination